A circular disk is cut from the following square sheet of metal. If each side of the square is 40 cm in length, what is the are of scrap metal when the circular disk is removed? Round your answer to the nearest square centimetre.
Answer by Cromlix(4381)

I am assuming that the circle's diameter is
equal to the side length of the square, namely 40cm.
Area of Square = length * width = 40*40 = 1600cm^2
Area of Circle = Pi*r^2 = Pi * 20^2 = 1256.6 cm^2
1600 - 1256.6 cm^2 = 343.4ins^2 left over.
Hope this helps.
:-)